1. if the voltage across an element is 8V and the current i entering the positive terminal is as shown, find the power delivered to the element at t=7 ms and the total charge and total energy delivered to the element between 1 and 10 ms.
  2. a copper wire has a resistance of 2 ohms it was drawn under pressure so that its diameter is reduced to one half its original diameter. the new resistance of wire is...
1
Expert's answer
2021-08-26T01:05:29-0400

Question Two:

 

The copper wire initially had: 2 Ohm of resistance

We can calculate this by:

We can let the original diameter to be d cm

The new diameter is reduced by 1.5

There the new d1  becomes: d/1.5 cm

Then the resistance of the wire after stretching is given by;

l×(rho)]/π r'² = 2 Ohm

 [2 × (rho)]/π r² = X Ohm

 Dividing d1 and d2, we get;

We get the new resistance by

(l/X) (d/d1)² = 1.5/2 = 0.75

Now in stretching the volume of the wire remains unchanged.

Initial volume of wire = π d² ×0.75 cm³

Extent of stretch = π d'² l cm

==> π d² ×X = π × d²× l ===> (d/d1')² = l.5/2

Inputting these figures for (d/d')² = l/X above, we obtain

(l/X) (l/X) = 9; ==> l² = 2×3×5.333=32 cm.

                              The new resistance becomes= 32Ohm